### v2

# Self-Calibrated CLIP for Training-Free Open-Vocabulary Segmentation

Abstract
Recent advancements in pre-trained vision-language models like CLIP, have enabled the task of open-vocabulary segmentation. CLIP demonstrates impressive zero-shot capabilities in various downstream tasks that require holistic image understanding. However, due to its image-level pre-training, CLIP struggles to capture local details, resulting in poor performance in segmentation tasks. Our analysis reveals that anomaly tokens emerge during the forward pass, drawing excessive attention from normal patch tokens, thereby diminishing spatial awareness. To address this issue, we propose Self-Calibrated CLIP (SC-CLIP), a training-free method that calibrates CLIP to produce finer representations while preserving its original generalization ability, without introducing new parameters or relying on additional backbones. Specifically, we first identify and resolve the anomaly tokens to mitigate their negative impact. Next, we enhance feature discriminability and attention correlation by leveraging the semantic consistency found in CLIP’s intermediate features. Furthermore, we explore how to effectively employ multi-level feature fusion under the training-free setting. Collectively, these strategies enhance CLIP’s feature representation with greater granularity and coherence. Experimental results demonstrate the effectiveness of SC-CLIP, achieving state-of-the-art results across all datasets and surpassing previous methods by 9.5%. Notably, SC-CLIP boosts the performance of vanilla CLIP ViT-L/14 by 6.8 times. Our source code is available here.
1 Introduction
Open-Vocabulary Segmentation (OVS) is an emerging task in computer vision that aims to segment arbitrary categories based on the textual inputs, overcoming the limitations of predefined category sets. To achieve this, models must generalize beyond the training data. Vision-language pretrained models such as CLIP [41], demonstrate remarkable zero-shot capabilities by leveraging large-scale image-text pairs, effectively fulfilling these requirements. However, CLIP’s image-text matching strategy leads to an excessive emphasis on global context, neglecting local and fine-grained details essential for dense prediction tasks. Consequently, directly applying CLIP to segmentation tasks yields poor performance. For example, as shown in Figure 1, the segmentation result generated by patch-text cosine similarity exhibits considerable noise. CLIP ViT-B/16 achieves only 8.9% mIoU on the COCO-Object dataset [6], significantly lagging behind its ability on image-level recognition.
Figure 1: Left: Vanilla CLIP produces a noisy segmentation map, while our Self-Calibrated CLIP (SC-CLIP) generates a clearer and finer result. Right: Comparison of the open-vocabulary segmentation performance, where SC-CLIP achieves the best results.
To address CLIP’s limitations in capturing local details, recent studies have proposed various modifications to its last layer. One line of research [74, 49, 4, 29, 25, 18, 22, 44] introduces correlative attention, replacing the original Q-K attention with alternatives like K-K, to enhance focus on relevant regions. But these methods still operate on the global and noisy inputs, hindering their effectiveness. Another approach [26, 54] incorporates additional backbones [7, 24, 43] to provide richer spatial details. Despite performance gains, they fail to fully exploit CLIP’s semantic knowledge and impose extra computational costs. In fact, both strategies overlook the underlying causes of CLIP’s global focus, preventing them from fundamentally resolving the issue of diminished spatial awareness in CLIP.
Motivated by this, we begin with an in-depth analysis of CLIP. As indicated by the orange circles in Figure 2 (a), we observe that different patch tokens consistently exhibit high activation regions within their attention map. These regions attract excessive attention from other normal patches, distracting their focus away from local and relevant areas. To further investigate, we apply PCA [1] analysis to CLIP’s features in Figure 2 (b), revealing these over-attended tokens differ significantly from the normal ones (including the [CLS] token). Thus, we refer to them as anomaly tokens. We attribute CLIP’s disadvantage in dense prediction tasks to the emergence of anomaly tokens, which leads to uniform attention activations across locations. This disrupts the attention’s ideal ability to extract relevant semantics, resulting in feature homogenization that diminishes local awareness and further exacerbates noise in the feature maps.
Figure 2: Anomaly tokens in CLIP. (a) visualizes the attention map of various selected patches (marked by $\bigstar$), all exhibit excessive focus on the same regions (indicated by orange circle), aligning with the identified outliers in the PCA analysis in (b).
Building on the analysis, we propose enhancing CLIP’s feature representation by weakening the influence of anomaly tokens. To this end, we introduce Self-Calibrated CLIP (SC-CLIP), a training-free approach that leverages CLIP’s inherent properties for effective calibration, strengthening its perception on local and relevant regions. On the one hand, we propose directly resolving these anomaly tokens. To identify them, we apply the Local Outlier Factor (LOF) algorithm [5], a method for detecting outliers. Once located, we replace these anomaly tokens with values interpolated from their spatial neighbors, considering that spatially close regions often share similar semantics. This not only serves as a regularization to prevent inappropriate attention focus but also reassigns meaningful semantic information to the anomaly tokens, aligning them with the local context. On the other hand, to relieve the feature homogenization problem caused by anomaly tokens, we propose a self-adjusting strategy to enhance feature discriminability and attention correlation. We observe that the intermediate layers’ features of CLIP exhibit good semantic coherence, as shown in Figure 4. In order to retain the rich semantics of deep features and the spatial consistency of intermediate features, we utilize the latter to adaptively aggregate deep features, while simultaneously enhancing attention correlation. This self-adjusting approach improves the overall semantic coherence. Furthermore, we explore how to effectively leverage multi-level feature fusion under the training-free setting and propose a two-pass strategy to enhance the capture of details at different scales.
Experimental results demonstrate that SC-CLIP achieves remarkable performance, establishing new state-of-the-art results across eight datasets, as shown in Figure 1. Our approach significantly outperforms previous methods by 9.5% on CLIP ViT-B/16. Notably, SC-CLIP boosts the performance of vanilla CLIP ViT-L/14 by 6.8 times, without the need for additional parameters, data, or backbones.
Our contributions can be summarized as follows:
  • We propose SC-CLIP, a training-free method designed to enhance CLIP’s dense feature representation, effectively addressing the uniform attention activations and feature homogenization caused by the anomaly tokens.
  • We calibrate CLIP with its inherent properties. Specifically, we propose resolving the anomaly tokens to mitigate their negative impact. And we utilize the spatial consistency in CLIP’s mid-level features to enhance feature discriminability and attention correlation.
  • Our approach sets new state-of-the-art results across popular benchmarks. And we conduct extensive experiments to validate the effectiveness of our method.
Figure 3: Resolving the Anomaly Tokens. (a) Illustration of the resolving process. We plot the feature map using the mean value of each token. After locating the anomaly tokens (the center of red square), we replace them with the interpolated values obtained from their neighboring regions. (b) Effect on the attention map. We highlight the changes for a normal token ($\bigstar$), and an anomaly token ($\blacktriangle$).
2 Related Work
Vision-Language Pre-training.
Vision-language models [21, 41, 67, 28, 70, 58] pretrained on large-scale web data, represented by CLIP [41], employ contrastive learning to align images with associated captions, demonstrating remarkable zero-shot capabilities across various downstream tasks that require holistic image understanding [23, 37, 2]. However, CLIP’s image-level pre-training, which relies on the [CLS] token to represent the whole image, causes the model to overly focus on global features at the expense of local and fine-grained details. This limitation hinders its performance in dense prediction tasks which require pixel-level understanding. To address this, our work focuses on effectively adapting CLIP for segmentation task while preserving its original knowledge and cross-modal capabilities.
Open-Vocabulary Segmentation.
The open-vocabulary segmentation (OVS) task focuses on segmenting images with arbitrary text queries by leveraging the zero-shot capabilities of vision-language models [21, 41]. This task has broad potential applications across many vision-related tasks [69, 35, 52, 51, 66, 65, 75, 72]. Existing works can be broadly classified into three categories: fully-supervised, weakly-supervised and training-free. Fully-supervised methods [16, 62, 13, 31, 61, 34, 63, 19, 68, 56, 27, 9, 64, 40] require fine-tuning on pixel-level annotated datasets. Weakly-supervised approaches [39, 42, 8, 59, 36, 60, 71, 20, 39] reduce reliance on dense annotations by using image-text pair to guide region grouping. Training-free methods [74, 49, 4, 29, 26, 54, 22, 44, 18, 3, 53, 33, 47] directly use CLIP for segmentation by making minimal adjustments to the model’s architecture without additional training.
Our approach needs no training and falls into the third category. Recent methods have discovered that CLIP’s final-layer exhibits poor spatial consistency and proposed various modifications. For instance, SCLIP [49] replaces the original Q-K attention with the combination of Q-Q and K-K attention, to enhance correlation. GEM [4] proposes generalized self-self attention and a set of regularizations. ClearCLIP [25] identifies the primary source of noise stems from the residual connections, and proposes to remove it. CLIPTrase [44] proposes using self-self attention along with clustering for post-processing. Other works leverage vision models like DINO [7] and SAM [24] for providing details. For example, CLIP-DINOiser [54] refines feature maps using the affinity learned from DINO’s feature correspondence, while ProxyCLIP [26] applies them to adjust attention weights. In contrast, our approach improves semantic coherence by exploiting CLIP’s internal properties and explicitly resolving the anomaly tokens.
3 Method
In this section, we begin with an overview of CLIP and its dense inference pipeline. We then present SC-CLIP, our approach aimed to enhance CLIP’s representation. First, we identify and mitigate the impact of anomaly tokens. Next, we utilize the mid-level features with spatial consistency to adaptively aggregate the deep features, while simultaneously enhancing attention correlation. Besides, we explore effective multi-level feature fusion to enrich details.
3.1 Preliminaries
CLIP ViT model encodes an input image into a token sequence $\mathbb{X}=[x_{\text{cls}},x_{1},\dots,x_{N}]$, where $x_{\text{cls}}$ is the [CLS] token and the others represent dense visual features, comprising $N$ patch tokens. The CLIP model includes multiple layers, with each layer $l$ processing the input $\mathbb{X}^{(l-1)}$ as follows:
$\mathbb{Z}^{l}$ $=\text{SA}(\text{LN}(\mathbb{X}^{(l-1)}))+\mathbb{X}^{(l-1)}$ (1)
$\mathbb{X}^{l}$ $=\text{FFN}(\text{LN}(\mathbb{Z}^{l}))+\mathbb{Z}^{l}$ (2)
where SA, FFN and LN denote self-attention module, feed-forward network, and layer normalization, respectively.
For dense inference, the visual features are aligned with $C$ categories to produce the patch-text similarity map of dimensions $N\times C$. And the final segmentation result is obtained by applying argmax operation to this similarity map.
Our method is training-free and modifies only the last layer of the CLIP visual encoder while keeping the other layers unchanged to prevent model collapse. For clarity and consistency in notation, we denote $\mathbb{X}^{penul}$ and $\mathbb{X}^{last}$ as the penultimate and last feature representations, respectively.
3.2 Resolving the Anomaly Tokens
We attribute CLIP’s limitations in dense prediction tasks to the presence of anomaly tokens within its features, which deviate significantly from normal tokens. These anomaly tokens cause other tokens to disproportionately focus on them in deep layers, leading to identical attention activation, which undermines attention’s capacity to extract semantically coherent regions. Since attention mechanism [48, 14] manages spatial arrangements, this pattern intensifies noise in the feature map, ultimately degrading performance. Studies [12, 55, 44] suggest that pretrained models like CLIP may identify redundant tokens and use them to gather global information, thereby expediting processing. However, these tokens lack semantics, conveying minimal information about their original positions.
To address the negative impact of anomaly tokens, we propose an intuitive approach to resolve them in $\mathbb{X}^{penul}$ before the last layer, as illustrated in Figure 3 (a). First, it is essential to identify the anomaly tokens. As previously analyzed in Figure 2, they exhibit a clear distinction from other tokens. To detect these anomalies, we employ the Local Outlier Factor (LOF) algorithm [5], a widely used method for anomaly detection. LOF identifies outliers by measuring the local density deviation of a data point relative to its neighbors. Specifically, it computes the local reachability density and assigns a high LOF score to points with significantly lower density than their neighbors, indicating potential anomalies. Besides, we implement the PyTorch-based LOF algorithm to enhance the computational efficiency.
Once the anomalies are located, we replace them with the values interpolated from their $3\times 3$ neighboring regions, based on the assumption that features in spatially adjacent regions are generally similar. Specifically, we apply a $3\times 3$ convolution kernel with the center set to 0. And if neighboring regions contain any other anomaly tokens, they are excluded from the interpolation. The operation’s formula is provided below, where $\tilde{\mathbb{X}}^{penul}$ denotes the feature after operation and $\mathcal{A}$ is the set of anomaly tokens.
$\tilde{\mathbb{X}}_{(x,y)}^{penul}$ $=\frac{\sum_{i=-1}^{1}\sum_{j=-1}^{1}w_{i,j}\cdot\mathbb{X}_{(x+i,y+j)}^{penul}}{\sum_{i=-1}^{1}\sum_{j=-1}w_{i,j}},\quad\forall(x,y)\in\mathcal{A}$
$w_{i,j}=\begin{cases}0,&\text{if }(x+i,y+j)\in\mathcal{A}\\ 1,&\text{otherwise}\end{cases}$ (3)
We believe that eliminating anomaly tokens offers two key benefits. First, it acts as a form of regularization for the attention. Second, it reassigns semantic information to these anomaly tokens, aligning them with the local context. As shown in the Figure 3 (b), the vanilla attention in CLIP has two issues: 1) normal tokens excessively focus on anomaly tokens, reducing attention on relevant local areas, and 2) anomaly tokens focus on prominent objects and other anomaly tokens. After resolving the anomaly tokens via interpolation, normal tokens refocus on relevant local regions, while the updated anomaly tokens also shift attention to appropriate areas. In summary, our method reduces the impact of anomaly tokens and enhances attention correlation.
3.3 Self-Adjusting for Semantic Coherence
After resolving the anomaly tokens, the model reduces focus on them in the last layer’s attention. However, a challenge still remains: anomaly tokens have already caused substantial disruption to the other normal tokens in the previous layers, diminishing their local awareness.
To enhance feature discriminability and correlation, some methods [26, 54] leverage backbones with strong spatial coherence like DINO [7], to provide fine-grained details. Although this approach improves performance, it introduces an additional backbone and incurs extra computational costs during either the training or testing phase.
Figure 4: Top: Visualization of patch similarities shows CLIP’s deep layers perform poorly, but its mid layers exhibit semantic consistency comparable to DINO. Bottom Left: ROC curve analysis further supports this, with SC-CLIP showing superior semantic coherence. Bottom Right: Detailed ROC analysis of our method.
Figure 5: Illustration of the self-adjusting strategy. (a) We use the similarity map from CLIP’s mid layer to adaptively aggregate deep features by combining semantically similar patches, resulting in a clearer segmentation map. The second row provides a detailed view of the process for the selected patch $\bigstar$. (b) We apply the similarity map to enhance attention, broadening and refining the activation regions.
This motivates us to explore whether such semantic coherence can be uncovered within CLIP itself. We begin by visualizing patch similarities, as shown in Figure 4 (top), where we observe the CLIP’s last layer performs poorly as expected. However, its mid-layer exhibits stronger semantic coherence, comparable to that of DINO. To further validate this observation, we quantitatively evaluate all features of CLIP using ROC curve analysis. Specifically, using the ADE20K dataset [73], we extract patches from each layer $l$ and compute cosine similarity between them to create the similarity map $\mathbf{Simi}^{l}\in\mathbb{R}^{N\times N}$, defined as $\mathbf{Simi}^{l}=\frac{\mathbb{X}^{l}\cdot\mathbb{X}^{l}}{\|\mathbb{X}^{l}\|\|\mathbb{X}^{l}\|}$. This similarity map serves as a binary classifier to indicate whether two patches belong to the same category, with patches of the same category labeled as 1 and otherwise 0. Each patch’s category is determined by majority voting on pixel labels in the segmentation map. For comparison, we also evaluate DINO. A higher area under the curve (AUC) in ROC analysis reflects better semantic consistency. As shown in Figure 4 (bottom left), CLIP’s mid-layer features exhibit strong spatial coherence (AUC=0.76), closely matching DINO (AUC=0.77). However, CLIP’s last-layer features perform much worse (AUC=0.66).
The OVS task requires spatially coherent cross-modal alignment. CLIP’s last feature $\mathbb{X}^{last}$ offers rich semantics but lacks coherence, while its mid-level features $\mathbb{X}^{mid}$ exhibit strong spatial consistency but are semantically limited. Inspired by [54], we propose feature aggregation to effectively combine the strengths of both. As illustrated in Figure 5 (a), we leverage $\mathbf{Simi}^{mid}$ to adaptively aggregate $\mathbb{X}^{deep}$, generating new features $\hat{\mathbb{X}}^{deep}$ by combining semantically similar patches with weighted contributions based on their similarity. This process can be formulated as follows:
$\hat{\mathbb{X}}_{p}^{deep}=\sum_{q=1}^{N}\text{Norm}(\mathbf{Simi}_{(p,\;q)}^{mid})\cdot\mathbb{X}_{q}^{deep}$ (4)
The aggregated features $\hat{\mathbb{X}}^{deep}$ show a clearer and more precise segmentation results compared to the original $\mathbb{X}^{deep}$.
Moreover, we find that the self-self attention (e.g., K-K attention) proposed in the previous works [25, 49, 29, 4] exhibit insufficient attention activations. Therefore, we further incorporate the similarity map $\mathbf{Simi}^{mid}$ to augment the attention operation, as detailed in the following formula. As shown in Figure 5 (b), the attention regions now display more extensive and accurate activations.
[…]
Feature aggregation and attention enhancement constitute the self-adjusting strategy, which significantly improves the semantic coherence, increasing the AUC to 0.80, as shown in Figure 4 (bottom right). This demonstrates that the inherent consistency within CLIP’s intermediate features can be effectively leveraged for adjustment.
3.4 Multi-level Feature Fusion
Motivated by the effectiveness of multi-level feature aggregation in enriching details for dense prediction tasks [32, 30, 57, 17, 29], we explore the potential of leveraging CLIP’s multi-level features in a training-free manner. A straightforward approach is to directly sum multi-level features: $\mathbb{X}^{last}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$, where $\mathcal{M}$ is the set of multi-level features. However, CLIP exhibits significant discrepancies across different layers. Specifically, the similarity between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is merely 0.094. Directly summing them severely disrupts CLIP’s cross-modal alignment capability. We identify two key principles for training-free feature fusion: 1) Ensuring the compatibility between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is crucial. This can be achieved by leveraging the parameter space of the last layer for alignment. The similarity between $\mathbb{X}^{last}$ and $\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right)$ increases to 0.983, where $\mathbf{L}$ denotes the last layer. This shows that using the last layer can significantly enhance feature compatibility. 2) Maintaining the integrity of $\mathbb{X}^{last}$, as it contains the direct alignment with the text embeddings.
Based on these insights, we propose a two-pass strategy, where CLIP’s final layer is explicitly employed for alignment. This strategy involves two forward passes—one with the original $\mathbb{X}^{penul}$ and another with multi-level features, formulated as: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right).$ Ablation experiments validate the effectiveness of our approach.
4 Experiments
4.1 Experiment Settings
Datasets and Metric We conduct comprehensive evaluations on eight commonly used benchmark datasets, which are grouped into two categories: 1) with a background class, including PASCAL VOC (VOC21) [15], PASCAL Context (Context) [38], and COCO Object (COCO-Obj) [6]; 2) without a background class, including PASCAL VOC20 (VOC20) [15], Cityscapes (City) [11], PASCAL Context59 (Context59) [38], ADE20K (ADE) [73], and COCOStuff (COCO-Stf) [6]. And we evaluate results with the standard mean-intersection-over-union (mIoU) as the metric.
Table 1: Performance comparison of our approach with other methods on eight semantic segmentation benchmarks. For a fair comparison, we reproduce all methods following the evaluation protocol in Section 4.1, considering the different settings used by each method. CLIP-DINOiser denotes our reproduced results using OpenAI’s pretrained weights. We report ProxyCLIP* results using its DINO-B/16 variant.
Method With a background category Without background category Avg.
VOC21 Context COCO-Obj VOC20 City Context59 ADE COCO-Stf
CLIP ViT-B/16
CLIP [41] 20.8 9.3 8.9 49.1 6.7 11.2 3.2 5.7 14.4
MaskCLIP [74] 51.4 22.5 24.9 62.9 25.6 26.2 12.3 16.9 30.3
ReCo [45] 25.1 19.9 15.7 57.7 21.6 22.3 11.2 14.8 23.5
GroupViT [59] 52.3 18.7 27.5 79.7 18.5 23.4 10.4 15.3 30.7
TCL [8] 51.2 24.3 30.4 77.5 23.5 30.3 14.9 19.6 33.9
CLIPSurgery [29] 55.2 30.3 29.7 77.5 33.1 33.4 16.1 22.2 37.2
LaVG [22] 62.1 31.6 34.2 82.5 26.2 34.7 15.8 23.2 38.8
NACLIP [18] 58.9 32.2 33.2 79.7 35.5 35.2 17.4 23.3 39.4
SCLIP [49] 59.7 31.7 33.5 81.5 32.3 34.5 16.5 22.7 39.1
ClearCLIP [25] 57.0 32.2 32.5 82.3 32.8 35.8 17.3 24.0 39.2
GEM [4] 58.7 32.0 32.9 81.7 32.6 35.6 16.9 23.9 39.3
CLIP-DINOiser [54] 57.8 32.1 32.0 78.0 36.2 36.0 16.8 22.6 38.9
ProxyCLIP* [26] 56.6 34.0 34.6 76.8 37.5 37.4 18.7 25.1 40.1
SC-CLIP (Ours) 64.6 36.8 37.7 84.3 41.0 40.1 20.1 26.6 43.9
CLIP ViT-L/14
CLIP [41] 10.3 4.5 4.4 19.9 3.2 5.7 1.9 3.2 6.6
MaskCLIP [74] 24.8 9.7 10.2 30.1 12.1 13.0 7.1 9.0 14.5
CLIPSurgery [29] 47.9 27.3 28.1 84.3 29.7 31.0 17.3 21.4 35.9
NACLIP [18] 52.1 28.7 29.9 78.6 31.4 32.1 17.3 21.4 36.4
SCLIP [49] 44.4 22.3 24.9 70.6 21.3 25.2 10.9 16.5 29.5
ClearCLIP [25] 48.6 28.0 28.6 84.8 32.1 31.5 16.9 21.2 36.5
GEM [4] 45.2 25.5 28.3 83.7 27.1 28.1 13.2 19.2 33.8
ProxyCLIP* [26] 58.1 34.1 37.4 82.0 38.1 37.3 21.2 25.5 41.7
SC-CLIP (Ours) 65.0 36.9 40.5 88.3 41.3 40.6 21.7 26.9 45.2
Implementation Details In our experiments, we utilize the CLIP [41] with ViT-B/16 and ViT-L/14 architectures. And our code implementation is built on the MMSegmentation [10]. For the evaluation protocol, we adopt the sliding window inference strategy from SCLIP [49]: input images are resized to have a short side of 336 (560 for Cityscapes [11] due to its higher resolution), and the slide inference is conducted with a $224\times 224$ window and a $112\times 112$ stride. No post-processing strategies are applied. Across all datasets, we use the standard ImageNet prompts [41] combined with their category names to construct text descriptions. For all the hyper-parameters, we keep consistent across all datasets without separate tuning.
4.2 Main Results
To ensure a fair comparison, we strictly follow the evaluation protocol specified in the Section 4.1 to reproduce all methods. For CLIP-DINOiser [54], we retrain the model using the OpenAI’s pretrained weights. Table 1 presents the comparison of all methods, where our SC-CLIP achieves an average mIoU of 43.9% on CLIP ViT-B/16, setting a new state-of-the-art results across eight benchmarks with a notable 9.5% improvement over previous methods. Furthermore, our method demonstrates robustness across different backbones, achieving optimal performance on CLIP ViT-L/14, improving the results by 3.5% average mIoU.
The vanilla CLIP achieves only a 14.4% and 6.6% mIoU on ViT-B/16 and ViT-L/14 respectively, demonstrating its limitations in capturing fine-grained spatial details. Our training-free method boosts its performance by threefold on the ViT-B/16 and by 6.8 times on the ViT-L/14.
In contrast to ProxyCLIP [26], which relies on the DINO-B/16 [7] backbone to provide attention weights, our method does not depend on any auxiliary backbone yet achieves superior results. This highlights the inherent properties of CLIP can be effectively leveraged to calibrate itself, thereby enhancing its feature representation.
4.3 Ablation Study
In this section, we conduct comprehensive ablation experiments to validate the effectiveness of our method. We use SCLIP [49] as our baseline, which modifies the attention mechanism in the final layer by replacing the original $\mathbf{Q}\mathbf{K}^{\top}$ attention with $\mathbf{Q}\mathbf{Q}^{\top}+\mathbf{K}\mathbf{K}^{\top}$ attention, to enhance correlations. Additionally, we remove the residual connections and feed-forward network (FFN) following [25, 26].
Table 2: Ablation experiments on the proposed strategies.
Method V21 C60 Obj C59 Stf Avg
baseline 58.2 32.3 34.0 35.4 23.6 36.7
+ AnomRes 60.3 33.3 34.9 36.5 24.3 37.9
+ AttnEnh 61.2 34.2 36.2 37.4 25.0 38.8
+ FeatAgg 62.1 34.8 37.4 38.0 25.7 39.6
+ MultiFuse 64.6 36.8 37.7 40.1 26.6 41.2
Analysis of Various Strategies
In Table 2, we incrementally incorporate each strategy to highlight its contribution. First, by resolving the anomaly tokens (denoted by AnomRes), our method achieves a 1.2% mIoU improvement, effectively addressing the issues brought by anomaly tokens, as discussed in Section 3.2. Next, we apply the self-adjusting strategy proposed in Section 3.3, which leverages the spatial consistency within CLIP’s mid-level features to enhance attention correlation (denoted by AttnEnh), leading to a 0.9% improvement. By adaptively aggregating deep features (FeatAgg), we obtain a further 0.8% gain in mIoU. Finally, the multi-level fusion strategy (MultiFuse) in Section 3.4 provides a 1.6% boost in mIoU. Collectively, these strategies contribute to a substantial 12.3% improvement over the baseline. Besides, we conduct a detailed ROC curve analysis as shown in Figure 4 (bottom right). SC-CLIP achieves an AUC of 0.81, surpassing the baseline by 17.4% and significantly enhancing the semantic coherence.
Table 3: Ablation study of resolving the anomaly tokens.
Method V21 C60 Obj C59 Stf Avg
baseline 58.2 32.3 34.0 35.4 23.6 36.7
resolve (1) 58.5 32.6 34.3 35.7 23.9 37.0
resolve (3) 59.2 32.9 34.6 36.0 24.0 37.3
resolve (5) 59.7 33.1 34.7 36.2 24.2 37.6
resolve (10) 60.3 33.3 34.9 36.5 24.3 37.9
resolve (15) 60.3 33.4 34.9 36.6 24.3 37.9
Table 4: Ablation study of the attention enhancement.
Method V21 C60 Obj C59 Stf Avg
only $\mathbf{K}\mathbf{K}^{\top}$ 60.6 33.1 34.3 36.6 24.2 37.8
only $\mathbf{Simi}^{mid}$ 60.4 33.8 35.8 37.1 24.8 38.3
both 61.2 34.2 36.2 37.4 25.0 38.8
Resolving the Anomaly Tokens
As we use the LOF algorithm to identify the anomaly tokens, we need to adjust the contamination hyperparameter to control the number of detected anomalies. We explore the optimal number of tokens to resolve, as shown in Table 3. As the number increases from 1 to 10, performance gradually improves. However, further increasing this number to 15 yields no additional gains. Therefore, we adopt the removal of 10 tokens (about 5% of the ViT-B/16 token sequence) in our method.
Self-adjusting Strategy
Table 4 shows the ablation study on attention enhancement. When both self-self attention ($\mathbf{K}\mathbf{K}^{\top}$) and $\mathbf{Simi}^{mid}$ are applied (Eq. (5)), the performance improves compared to using either component alone, demonstrating the effectiveness of attention enhancement. And in Table 5, we conduct a comprehensive analysis of the feature aggregation. We explore different sources to provide similarity map. And we find that adjusting both $\mathbb{X}^{penul}$ and $\mathbb{X}^{last}$ is essential, serving as the pre- and post-adjustment for the last layer. Experiments indicate that using layer 9 for pre-adjustment and layer 4 for post-adjustment yields the optimal results, corresponding to (9, 4) in the table, achieving an average 39.8% mIoU, a 6% improvement over the baseline. We hypothesize that deeper mid-layers (e.g., 9), with their broader activations, aggregate richer semantics that help the attention mechanism capture more coherent regions. In contrast, shallower mid-layer (e.g., 4), with more localized activations, can preserve intricate spatial information. And our method is robust to different combinations, with (8, 5) also achieving an mIoU of 39.5. More results are available in the supplementary materials. Furthermore, we compare our approach with DINO, following prior studies [54, 46, 50]. When applied to $\mathbb{X}^{last}$, DINO shows superior performance, showing its fine-grained details; however, our approach achieves comparable results when both pre- and post-adjustments are applied, fully unlocking CLIP’s potential without relying on additional backbones.
Table 5: Ablation study of the feature aggregation.
Position Source V21 C60 C59 Stf Avg
without 58.2 32.3 35.4 23.6 37.4
$\mathbb{X}^{penul}$ DINO 58.0 31.8 34.3 23.5 36.9
4 58.9 32.6 35.2 24.2 37.7
8 60.1 33.5 36.2 24.5 38.6
9 60.0 33.4 36.2 24.5 38.5
$\mathbb{X}^{last}$ DINO 62.6 34.0 37.2 25.4 39.8
4 61.4 34.2 37.2 25.1 39.5
5 61.2 34.1 37.1 25.0 39.4
9 60.1 33.7 36.7 24.7 38.8
both DINO 58.8 32.0 34.5 23.8 37.3
(4, 9) 58.5 32.5 35.1 24.1 37.6
(8, 5) 61.7 34.2 37.1 25.1 39.5
(8, 4) 61.9 34.3 37.1 25.2 39.6
(9, 4) 62.0 34.4 37.3 25.3 39.8
Table 6: Ablation study of the multi-level fusion strategy.
Method V21 C60 Obj C59 Stf Avg
dual path [29] 63.5 35.4 37.3 38.5 26.2 40.2
①baseline 62.1 34.8 37.4 38.0 25.7 39.6
②direct sum 19.5 3.5 2.7 15.4 8.0 9.8
③one pass 63.6 36.1 35.2 39.7 26.1 40.1
two pass 64.6 36.8 37.7 40.1 26.6 41.2
Table 7: Selected layers of the multi-level fusion strategy.
Layers w/o 1-10 2-10 3-10 4-10 5-10 4-9 4-11
Avg 42.4 43.8 43.9 43.9 43.9 43.8 43.8 43.6
Figure 6: Qualitative Results of Open-Vocabulary Segmentation. We compare our method with CLIP [41], MaskCLIP [74], SCLIP [49] and ClearCLIP [25], all without post-processing. Our SC-CLIP produces much clearer and more accurate segmentation maps.
Multi-level Fusion Strategy
We conduct ablation study of the multi-level fusion strategy in Table 6 with following approaches. ①baseline: $\mathbf{L}(\mathbb{X}^{penul})$direct sum: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$one pass: $\mathbf{L}(\mathbb{X}^{penul}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i})$two pass: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right)$. Here, $\mathbf{L}$ is the last layer, and $\mathcal{M}$ is the multi-level features. The results align with our analysis in Section 3.4, validating two key principles: (1) using the last layer for alignment: comparison between ② and ④ reveals that ④ aligns the multi-level features using the parameter space of the last layer, which improves feature compatibility and leads to a significant performance boost. (2) maintaining the integrity of the last feature: comparison between ③ and ④ indicates that ③ weakens the original last feature, resulting in a performance drop. Additionally, we compare our approach with the dual-path strategy proposed in CLIPSurgery [29], where each layer of the image encoder in the new path is modified to self-self attention, and the outputs of the new path are directly aggregated. This can be formulated as: dual path: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}_{new}$, where $\mathbb{X}^{i}_{new}$ is the feature from the new path. This strategy can be regarded as an enhanced version of ②. However, in comparison, our method achieves superior results while avoiding the additional computational overhead introduced by the new path.
To analyze the impact of selected layers in multi-level fusion, we report the average mIoU across eight datasets in Table 7, indicating that our method is robust to the choice of selected layers, as different combinations yield similar results. The notation 4–10 in the table denotes fusing all features from layer 4 to 10, which achieves the best 43.9% mIoU, surpassing the baseline (w/o) by 1.5% mIoU.
Efficiency Comparison
As shown in Table 8, we present the efficiency comparison on the COCO-Object dataset using a V100 GPU. Compared to the previous state-of-the-art method, ProxyCLIP [26], SC-CLIP requires no additional backbones, significantly improving efficiency. Specifically, the FPS increase from 3.9 to 6.5, and the FLOPs are reduced from 34.4G to 17.5G. Additionally, we show the detailed breakdown of the computational cost for each component.
Table 8: Efficiency comparison of training-free methods.
Models FLOPs(G) $\downarrow$ Params(M) $\downarrow$ Speed(FPS) $\uparrow$
CLIP 17.7 149.6 7.5
ProxyCLIP 34.4 235.4 3.9
Ours
baseline 16.8 149.6 7.8
+LOF 16.8 149.6 7.6
+AnomRes 16.8 149.6 7.1
+SelfAdjust 16.9 149.6 6.7
+MultiFuse 17.5 149.6 6.5
5 Visualization
In Figure 6, we present the qualitative results of various methods, showing that SC-CLIP consistently achieves more precise and higher quality results than other methods. In contrast, CLIP’s results are considerably noisy and exhibit a homogeneity effect (e.g., in the 2nd row, the entire image is labeled as “dog”). By performing self-calibration, SC-CLIP enables CLIP to produce semantically coherent results.
6 Conclusion
We present Self-Calibrated CLIP (SC-CLIP), a training-free method to enhance the open-vocabulary segmentation performance of CLIP. We first identify and resolve the negative impact of anomaly tokens. Then we leverage the semantic consistency within CLIP’s intermediate features to enhance both feature discriminability and attention correlation. Together with multi-level fusion strategy, SC-CLIP sets new state-of-the-art results, without requiring additional data, parameters or backbone. This demonstrates the inherent properties of CLIP can be effectively leveraged to calibrate itself to produce semantically coherent representations.